# PyMieDiff: A differentiable Mie scattering library

### v2

@@ Abstract
Light scattering by spherical-shaped particles of sizes comparable to the wavelength is foundational in many areas of science, from chemistry to atmospheric science, photonics and nanotechnology. With the new capabilities offered by machine learning, there is a great interest in end-to-end differentiable frameworks for scattering calculations. Here we introduce PyMieDiff, a fully differentiable, GPU-compatible implementation of Mie scattering for core-shell particles in PyTorch. The library provides native, autograd-compatible spherical Bessel and Hankel functions, vectorized evaluation of Mie coefficients, and APIs for computing efficiencies, angular scattering, and near-fields. All inputs – geometry, material dispersion, wavelengths, and observation angles and positions – are represented as tensors, enabling seamless integration with gradient-based optimisation or physics-informed neural networks. The toolkit can also be combined with “TorchGDM” for end-to-end differentiable multi-particle scattering simulations. PyMieDiff is available under an open source licence at https://github.com/UoS-Integrated-Nanophotonics-group/MieDiff.
@@ I Introduction
A variety of open-source packages implement Mie theory for spheres. For example, “MiePython” is a pure-Python (NumPy/Numba) implementation of Mie theory for homogeneous spheres.Prahl (2025) “PyMieSim” is an open-source Python/C++ toolkit that supports scattering from spheres, infinite cylinders and core–shell geometries.de Sivry-Houle et al. (2023) “Scattnlay” is a c++ programm, focused on the calculation of nearfields inside and around multi-layer spherical particles.Ladutenko et al. (2017) “pyMieCS” is a vectorized NumPy library specialized for core–shell nanoparticles.Wiecha (2024) These tools offer validated, high-speed forward solvers, but none provide native differentiation capabilities or GPU backends.
@@
Automatic differentiation (AD) is the key numerical technique behind deep learning, and allows to calculate arbitrary derivatives for any numerical calculation with close to analytical precision.Wengert (1964) The photonics community has seen rapid growth of differentiable simulation tools based on AD. Recent efforts have produced open-source Maxwell solvers with autodifferentiation (e.g., FEM, Fourier modal, FDTD or volume integral approaches) to enable gradient-based inverse design and topology optimization.Hughes et al. (2019); Minkov et al. (2020); Colburn and Majumdar (2021); Wang et al. (2022); Luce et al. (2024); Kim et al. (2024); Ponomareva et al. (2025) For Mie computations however, so far still external libraries are required, which are not AD-compatible.Ponomareva et al. (2025)
@@
Recent works have highlighted the power of integrating physics-based models with deep learning. An important example are deep learning based nanoparticle design approaches which, so far, often rely on data-based tandem networks or surrogate models.Liu et al. (2018); So et al. (2019); Estrada-Real et al. (2022); Khaireh-Walieh et al. (2023) These solutions often suffer from a lack of fidely of the surrogate models Dinsdale et al. (2021), as their interpolation is limited to the dataset they are trained on. These large datasets only cover a limited parameter space, and new problems reqiure new datasets. Replacing such surrogate models by AD capable analytical solvers would suppress this typical point of failure.
@@
We present the PyTorch-based toolkit “PyMieDiff” which implements analytical Mie theory for core–shell spherical particles with full support for automatic differentiation. Our library implements the standard Mie recurrence (spherical Bessel/Hankel forward and backward recurrences, angular functions, vector spherical harmonicsCai (2011)) entirely in PyTorch. All quantities – particle geometry (layer thicknesses, materials), wavelengths, and scattering angles – can be treated as differentiable tensors. As a result any Mie-derived observable can be backpropagated to compute gradients with respect to any input parameter. The code is fully vectorized over Mie orders, wavelengths, angles, and positions. It supports batched evaluation of many particles in parallel and runs on GPU. Our simple API exposes a Particle class for defining core-shell spheres with possibly dispersive materials (using the refractiveindex.info formatPolyanskiy (2024)). It implements methods to compute extinction, absorption and scattering efficiencies, far-fields and S-matrix elements, as well as near-fields, all with autograd support, as illustrated in figure 1.
@@
  • Flexible materials interface: refractive index data (from refractiveindex.info) are interpolated through a PyTorch implementation, so material dispersion enters the gradient graph transparently.
@@
Figure 1: Overview of PyMieDiff’s features. The core of the code is a differentiable Mie solver, which is used to compute various differentiable observables (efficiencies, angular scattering, near-fields). All calculations are fully implemented in PyTorch, they all support automatic differentiation and run on GPU. The shown results are from a gold-silicon core-shell particle with core radius $r_{c}=100\,$nm and shell radius $r_{s}=200\,$nm, placed in vacuum and illuminated by a linearly polarized plane wave.
@@ II Implementation
Mie theory describes the interaction of light with a spherical particle through field expansion coefficients $a_{n},b_{n},c_{n}$ and $d_{n}$, where $a_{n}$ and $b_{n}$ are the coefficients for the outgoing (scattered) field, while $c_{n}$ and $d_{n}$ are the expansion coefficients for the incoming (internal) fields.
In the case of a core-shell particle of core radius $r_{c}$ and shell radius $r_{s}$, the Mie coefficients are functions of the scale factors $x=kr_{c}$ and $y=kr_{s}$ (wavevector $k=k_{0}n_{\text{env}}$, $k_{0}=2\pi/\lambda_{0}$) and the relative refractive indices, $m_{c}=n_{c}/n_{\text{env}}$, and $m_{s}=n_{s}/n_{\text{env}}$ for the core and shell, respectively. The explicit form of the core-shell Mie scattering coefficients is given in the appendix .6. The evaluation of the Mie coefficients requires Riccati-Bessel functions and their derivatives:
$\psi_{n}(z)=zj_{n}(z),\quad\xi_{n}(z)=zh_{n}^{(1)}(z),\quad\chi_{n}(z)=zy_{n}(z).$ (1)
@@
The derivatives can be obtained through the identity:
$\frac{d}{dz}f_{n}(z)\;=\;f_{n-1}(z)-\frac{n+1}{z}\,f_{n}(z)\,$, (2)
where $f$ represents any spherical Bessel or Hankel function (see also appendix .4).
In PyTorch (as of version 2.9), spherical Bessel functions are not implemented. The key contribution of our work is therefore to provide PyTorch-based, AD-capable, vectorized, fast and stable spherical Bessel routines. We implement two versions of AD enabled spherical Bessel functions: (1) a SciPy wrapper and (2) a native PyTorch implementation based on recurrences. While the SciPy wrapper is based on stable and tested SciPy routines, it is not GPU compatible and comes with additional memory transfer overhead. This limits its efficiency, especially in batched evaluations. The native torch implementation on the other hand is fully GPU capable and entirely vectorized. On single, non-parallelized evaluations the SciPy version is faster, whereas in cases where many wavelengths, several particles, high orders, or multiple angles are to be calculated, the native PyTorch implementation becomes advantageous.
In our PyTorch implementation we compute the spherical Bessel functions $j_{n}$ using downward recurrence, while the spherical Neumann functions $y_{n}$ are obtained by upward recurrence from analytic low-order seeds. This separation is necessary because forward propagation of the $j_{n}$ recurrence rapidly becomes unstable for large $n$ and for $z$ with a large imaginary part (e.g., for spheres with strongly absorbing / metallic permittivities).
Figure 2: Benchmarks. (a) timing for increasing numbers of simultaneously evaluated wavelengths. (b) timing for increasing numbers of simultaneously evaluated particles. (c) parallelization speedup on multi-core CPUs. (d) timing comparison to other Mie toolkits (on CPU). The batched example evaluates 256 different particles simultaneously, each at 256 wavelengths. Note that “treams” is not performance-optimized for Mie evaluation since it is not a dedicated Mie toolkit (but T-matrix), in consequence the time-bar for the treams result is not fully shown.
Using the Riccati-Bessel form $\psi_{n}(z)=z\,j_{n}(z)$ the three-term recurrence reads:
$\psi_{n+1}(z)=\frac{2n+1}{z}\,\psi_{n}(z)-\psi_{n-1}(z)\,.$ (3)
To obtain $j_{n}$, we run through Eq. (3) as backward recurrence, starting from seeds $j_{n}=0$ and $j_{n+1}=10^{-25}$ at a sufficiently large orderCai (2011). The sequence is finally re-normalized to the known $j_{0}$ ($j_{0}(z)=\sin z/z$). The Neumann functions $y_{n}$ are generated via the same three-term recurrence, applied upward from analytic seeds ($y_{0}(z)=-\cos z/z$, $y_{1}(z)=-\cos z/z^{2}-\sin z/z$).Gillman and Fiebig (1988) Implementing both branches using only differentiable PyTorch tensor operations preserves automatic differentiation, while avoiding overflow/underflow and the loss of accuracy encountered with lossy materials, for example in plasmonic parameter regimes. Spherical Hankel functions and derivatives are generated from the spherical Bessel relations (see Eqs. (2) and appendix .4).
With these spherical Bessel and Hankel functions and their derivatives, the Mie coefficients can be readily evaluated (see appendix .6). Based on these, scattering efficiencies and angular scattering patterns can be calculated, as detailed in appendix .7.
@@
12 r_core=r_core,
13 r_shell=r_shell,
14 mat_core=mat_core,
15 mat_shell=mat_shell,
16 mat_env=n_env,
17)
18
19# --- evaluation
20wl0 = torch.linspace(500, 1000, 50) # nm
21k0 = 2 * torch.pi / wl0
@@
23# efficiency spectra
24cs = p.get_cross_sections(k0)
25
26# angular scattering (angles in radian)
27theta = torch.linspace(0.0, 2 * torch.pi, 100)
28angular = p.get_angular_scattering(k0, theta)
29
30# nearfields (positions in nm)
31x, z = torch.meshgrid(
32 torch.linspace(-250, 250, 100),
33 torch.linspace(-250, 250, 100),
34)
35y = torch.ones_like(x)
36r_probe = torch.stack([x, y, z], dim=-1)
37
38fields = p.get_nearfields(k0=k0, r_probe=r_probe)
@@
III Benchmark
In figure 2 we assess the computational performance of PyMieDiff and compare it against existing Mie scattering toolkits. Owing to its fully vectorized implementation utilizing PyTorch tensors, PyMieDiff exhibits efficient scaling for any independent parameter, like number of wavelengths per spectrum (Fig. 2a), incident angles, evaluation positions or the number of particles (Fig. 2b). The evaluation time saturates only, when the memory transfer overhead is of similar magnitude as the actual computation time of one calculation batch.
Note that, while GPU execution is supported, the performance is currently constrained by the recurrences operating internally at double precision, which is required for numerical stability. On consumer grade GPUs (for Fig. 2 an NVIDIA RTX 4090 was used), double precision is significantly less efficient than lower precision types, which explains the superior performance on CPUs. Yet, GPU support can still benefit integration into deep learning schemes, as memory transfer during model evaluation is reduced.
On multi-core CPUs, PyMieDiff demonstrates good parallel efficiency, which seems to approach a plateau after eight cores (Fig. 2c). These parallelization capabilities are automatically used since they are inherent to PyTorch. This scaling behaviour is particularly beneficial for large-scale parameter sweeps, many-particle optimization tasks or machine learning integration with batched training schemes. In a direct comparison with other publicly available toolkits (Fig. 2d), PyMieDiff achieves markedly lower evaluation times per wavelength as soon as batched evaluation is possible, when our toolkit provides more than one order-of-magnitude improvement over the other evaluated tools. Note that while the T-Matrix package “treams”Beutel et al. (2024) is included for completeness, it is not parallelized and not specifically optimized for single particle Mie calculations.
@@ III Benchmark
Figure 4: Optimization example. (a) Sketch of the particle optimization procedure. A large number of random core-shell particles is randomly initialized and evaluated using PyMieDiff. Their optical properties (orange lines, here scattering efficiency spectra) are compared with the design target (dotted black lines, here a Gaussian response). The derivatives of the loss function (here MSE) with respect to the geometry parameters are obtained via PyTorch autograd and used to iteratively update the geometries. (b) Sketches of the best particles of selected iteration during an optimization (top row) together with their spectral response (bottom row). Relative sizes are on the same scale, the shades of the colors indicate the real part of the refractive index (darker means higher). (c) Loss function during the optimization.
@@ IV.1 Forward evaluation
We compared the results to several other, openly available Mie solvers and onbtained results identical to machine precision. This comparison can be found in the examples of the online documentation.
Figure 5: “Mie-informed” tandem network. (a) Sketch of the training configuration. The design observable (e.g., scattering efficiency, or any other Mie-calculated value) is fed into the “inverse network” (here a simple MLP). The NN result is interpreted as core-shell geometry parameters and fed into PyMieDiff, which calculates analytically the response of the particle. The result of the Mie calculation is compared to the original input data using some loss function (here MSE). The inverse network is trained using backpropagation of the loss through the differentiable Mie solver and then the network itself. The inset shows the training loss for an inverse network trained on scattering efficiency spectra of dielectric particles. (b) Examples of inverse design responses (orange lines) vs. randomly sampled design targets (dashed black lines) from the training dataset .

### v3

@@ Abstract
Light scattering by spherical-shaped particles of sizes comparable to the wavelength is foundational in many areas of science, from chemistry to atmospheric science, photonics and nanotechnology. With the new capabilities offered by machine learning, there is a great interest in end-to-end differentiable frameworks for scattering calculations. Here we introduce PyMieDiff, a fully differentiable, GPU-compatible implementation of Mie scattering for layered, spherical particles in PyTorch. The library provides native, autograd-compatible spherical Bessel and Hankel functions, vectorized evaluation of Mie coefficients, and APIs for computing efficiencies, angular scattering, and near-fields. All inputs – geometry, material dispersion, wavelengths, and observation angles and positions – are represented as tensors, enabling seamless integration with gradient-based optimisation or physics-informed neural networks. The toolkit can also be combined with “TorchGDM” for end-to-end differentiable multi-particle scattering simulations. PyMieDiff is available under an open source licence at https://github.com/UoS-Integrated-Nanophotonics-group/MieDiff.
@@ I Introduction
A variety of open-source packages implement Mie theory for spheres. For example, “MiePython” is a pure-Python (NumPy/Numba) implementation of Mie theory for homogeneous spheres.Prahl (2025) “PyMieSim” is an open-source Python/C++ toolkit that supports scattering from spheres, infinite cylinders and core–shell geometries.de Sivry-Houle et al. (2023) “Scattnlay” is a c++ program, focused on the calculation of nearfields inside and around multi-layer spherical particles.Ladutenko et al. (2017) “pyMieCS” is a vectorized NumPy library specialized for core–shell nanoparticles.Wiecha (2024) These tools offer validated, high-speed forward solvers, but none provide native differentiation capabilities or GPU backends.
@@
Automatic differentiation (AD) is the key numerical technique behind deep learning, and allows calculating arbitrary derivatives for any numerical calculation with close to analytical precision.Wengert (1964) The photonics community has seen rapid growth of differentiable simulation tools based on AD. It is a generalization of the commonly used adjoint method.Capers et al. ; Grieshammer et al. (2024); Bahmani et al. (2025) While the adjoint method requires manual, analytical derivation of reciprocal physics to obtain gradients, AD programmatically tracks every operation within a generalized computational graph to enable backward calculation of arbitrary derivatives. AD provides high-level flexibility at the cost of a significant code complexity and memory overhead, required to store the computational graph. Recent efforts have produced open-source Maxwell solvers with automatic differentiation (e.g., FEM, Fourier modal, FDTD or volume integral approaches) which enable gradient-based inverse design and AD-driven topology optimization.Hughes et al. (2019); Minkov et al. (2020); Colburn and Majumdar (2021); Wang et al. (2022); Luce et al. (2024); Kim et al. (2024); Ponomareva et al. (2025); Mahlau et al. (2026) These tools have been implemented using a range of scripting languages and libraries. The majority of recent tools are written in Python, where currently the most commonly used autodiff libraries are PyTorch and JAX. For Mie theory computations however, no autodiff capable tool exists so far.
@@
Recent works have highlighted the power of integrating physics-based models with deep learning. An important example are deep learning based nanoparticle design approaches which, so far, often rely on data-based tandem networks or surrogate models.Liu et al. (2018); So et al. (2019); Estrada-Real et al. (2022); Khaireh-Walieh et al. (2023) These solutions often suffer from a lack of surrogate model fidelity Dinsdale et al. (2021), as their interpolation is limited to the dataset they are trained on. These large datasets only cover a limited parameter space, and new problems require new datasets. Replacing such surrogate models by AD capable analytical solvers would suppress this typical point of failure.
@@
We present the PyTorch-based toolkit “PyMieDiff” which implements analytical Mie theory for core–shell spherical particles with full support for automatic differentiation. Our library implements the standard Mie recurrence (spherical Bessel/Hankel forward and backward recurrences, angular functions, vector spherical harmonicsCai (2011)) entirely in PyTorch. All quantities – particle geometry (layer thicknesses, materials), wavelengths, and scattering angles – can be treated as differentiable tensors. As a result any Mie-derived observable can be backpropagated to compute gradients with respect to any input parameter. The code is fully vectorized over Mie orders, wavelengths, angles, and positions. It supports batched evaluation of many particles in parallel to maximize computational efficiency. Thanks to the native pytorch implementation it also runs on GPU.
Our simple API exposes a Particle class for defining core-shell spheres with possibly dispersive materials (using the refractiveindex.info formatPolyanskiy (2024)). It implements methods to compute extinction, absorption and scattering efficiencies, far-fields and S-matrix elements, as well as near-fields, all with autograd support, as illustrated in figure 1.
@@
  • Flexible materials interface: PyMieDiff supports using refractive index data for real materials (from refractiveindex.info jaml files). Dispersion data is interpolated through a PyTorch-based interpolation routine, so material dispersion enters the gradient graph transparently.
@@
Figure 1: Overview of PyMieDiff’s features. The core of the code is a differentiable Mie solver, which is used to compute various differentiable observables (efficiencies, angular scattering, near-fields). All calculations are fully implemented in PyTorch, they all support automatic differentiation and run on GPU. The shown results are from a gold-silicon-titania 3-layer particle with core radius $r_{1}=50\,$nm, inner layer radius $r_{2}=150\,$nm and outer shell radius $r_{2}=210\,$nm, placed in vacuum and illuminated by a linearly polarized plane wave. The shown spectrum goes from 500 nm to 1400 nm. The angular scattering is calculated at 770 nm and the near-fields at 525 nm. Material dispersion is interpolated from “refractiveindex.info”.
@@ II Implementation
Mie theory describes the interaction of light with a spherical particle through field expansion with spherical waves, using the coefficients $a_{n},b_{n},c_{n}$ and $d_{n}$, where $a_{n}$ and $b_{n}$ are the coefficients for the outgoing (scattered) fields, while $c_{n}$ and $d_{n}$ are the expansion coefficients for the incoming fields (required in the expansion of internal fields).
In the case of a layered particle, the Mie coefficients are functions of the scale factors $x_{l}=kr_{l}$ (with wavevector $k=k_{0}n_{\text{env}}$; $k_{0}=2\pi/\lambda_{0}$; and radius $r_{l}$ of the $l$th spherical layer) and the relative, complex refractive indices $m_{l}=n_{l}/n_{\text{env}}$ for layer $l$. The explicit form of the Mie scattering coefficients is given in the appendix .6. The evaluation of the Mie coefficients requires Riccati-Bessel functions and their derivatives:
$\psi_{n}(z)=zj_{n}(z),\quad\xi_{n}(z)=zh_{n}^{(1)}(z),\quad\chi_{n}(z)=-zy_{n}(z).$ (1)
@@
Implementing Mie series directly based on these spherical Bessel functions can become numerically unstable for large size parameters or absorbing materials.Wiscombe (1980) Therefore, modern implementations reformulate the Mie expansion using logarithmic derivatives of the form
$D_{n}^{(1)}(z)\;=\;\frac{\psi_{n}^{\prime}(z)}{\psi_{n}(z)}\,.$ (2)
These can be numerically calculated through similar recurrences as for the normal spherical Bessel functions. An important detail is that upward recurrences are stable for the logarithmic derivatives of $\xi_{n}$, while downward recurrences are needed to calculate the logarithmic derivatives for $\psi_{n}$.Yang (2003) All mathematical and algorithmic details can be found in the appendix .6. Our implementation closely follows the work of Peña and Pal.Peña and Pal (2009)
In PyTorch (as of version 2.10), neither spherical Bessel functions, nor their logarithmic derivatives are implemented. The key contribution of our work is therefore to provide PyTorch-based, AD-capable, vectorized, fast and stable spherical Bessel routines. We implement following AD enabled functions: (1) A PyTorch wrapper to the SciPy versions of standard spherical Bessel functions and their derivatives. (2) A native PyTorch implementation of spherical Bessel functions and their derivatives based on up- and down-recurrences. (3) A native PyTorch implementation of logarithmic derivatives and all further ingredients to implement the stable Mie algorithm, suggested by Yang and further improved by Peña and Pal.Yang (2003); Peña and Pal (2009) The spherical Bessel and Hankel functions were mainly implemented for testing. But since they may be useful for various other applications that require torch-versions of these functions with native GPU support, all special functions are available through the “pymiediff.special” module.
Based on the logarithmic derivatives, PyMieDiff implements a native PyTorch Mie theory stack comprising Mie coefficients, angular functions, and vector spherical harmonics. This is then used to implement functions that evaluate crossections, angular radiation, and near-fields.
@@
Figure 2: Benchmarks. If not otherwise noted, benchmarks were run for a single dielectric core-shell particle at 256 wavelengths. (a) timing for increasing numbers of simultaneously evaluated wavelengths. (b) timing for increasing numbers of simultaneously evaluated particles (single wavelength). (c) parallelization speedup on multi-core CPUs. (d) timing for increasing nr. of shell layers. (e) timing comparison to other Mie toolkits (on CPU). The cases “batched” and “GPU” simultaneously evaluate 256 different particles, at 256 wavelengths each. Our CPU is a Zen 3 AMD processor. Our GPU is an NVIDIA RTX 4090. Note that “treams” is not performance-optimized for Mie evaluation since it is not a dedicated Mie toolkit (but T-matrix), in consequence the time-bar for the treams result is not fully shown.
III Benchmark
In figure 2 we assess the computational performance of PyMieDiff and compare it against existing Mie scattering toolkits. Owing to its fully vectorized implementation utilizing PyTorch tensors, PyMieDiff exhibits efficient scaling for any independent parameter, like number of wavelengths per spectrum (Fig. 2a), scattering angles, evaluation positions, the number of particles (Fig. 2b). The number of layers requires evaluation of a cascade of interdependent scattering coefficients (see appendix .6), thus it scales linearly with compute (Fig. 2d). We observe that the evaluation time is dictated by memory transfer until around the order of $10^{3}$ parallel evaluations. Running PyMieDiff on GPU shows that the performance is memory-bound until at least $10^{5}$ concurrent Mie evaluations. GPU becomes advantageous only for large batch sizes above $\approx 10^{3}$ (see figures 2a-b). But even for lower batch sizes, GPU support can still benefit integration into deep learning schemes, as memory transfer during model evaluation is reduced and the neural network parts of the calculation may strongly benefit from GPU evaluation also on smaller batch sizes.
@@ III Benchmark
12 r_layers=[r_core, r_shell],
13 mat_layers=[mat_core, mat_shell],
14)
15
16# --- evaluation
17wl0 = torch.linspace(500, 1000, 50) # nm
18k0 = 2 * torch.pi / wl0
19
20# efficiency spectra
21cs = p.get_cross_sections(k0)
@@
23# angular scattering (angles in radian)
24theta = torch.linspace(0.0, 2 * torch.pi, 100)
25angular = p.get_angular_scattering(k0, theta)
26
27# nearfields (positions in nm)
28x, z = torch.meshgrid(
29 torch.linspace(-250, 250, 50),
30 torch.linspace(-250, 250, 50),
31)
32y = torch.ones_like(x)
33r_probe = torch.stack([x, y, z], dim=-1)
34
35fields = p.get_nearfields(k0=k0, r_probe=r_probe.view(-1, 3))
@@
On multi-core CPUs, PyMieDiff demonstrates good parallel efficiency, slowing down for eight or more cores (Fig. 2d). We attribute this reduction in scaling efficiency to memory transfer overhead in PyTorch’s shared-memory parallelization, not optimized for embarrassingly parallel tasks. We note that these parallelization capabilities are automatically used since they are inherent to PyTorch. This scaling behaviour is particularly beneficial for large parameter sweeps, many-particle optimization tasks or machine learning integration with batched training schemes. In a direct comparison with other publicly available toolkits (Fig. 2d), PyMieDiff is on par with the fastest implementation (scattnlay) and achieves markedly lower evaluation times per wavelength as soon as batched GPU evaluation is possible. At best, our toolkit provides more than one order-of-magnitude improvement over the fastest of the other evaluated tools. Note that while the T-Matrix package “treams”Beutel et al. (2024) is included for completeness, it is not parallelized and not specifically optimized for single particle Mie calculations.
@@
Figure 4: Example of gradient-based particle optimization. (a) Sketch of the particle optimization procedure. A large number of random core-shell particles is randomly initialized and evaluated using PyMieDiff. Their optical properties (orange lines, here: scattering efficiency spectra) 21 points between 400400\,nm and 800800\,nm are compared with the design target (dotted black lines, here: a Gaussian response). The derivatives of the loss function (here MSE) with respect to the geometry parameters are obtained via PyTorch autograd and used to iteratively update the geometries. (b) Sketches of the best particles of selected iteration during an optimization (top row) together with their spectral response (bottom row). Relative sizes are on the same scale, the shades of the colors indicate the real part of the refractive index (darker means higher). (c) Loss function during the optimization.
@@ IV.1 Forward evaluation
We compared the results to several other, openly available Mie solvers and obtained results identical to machine precision. This comparison can be found in the examples of the online documentation.
1import torch
2import pymiediff as pmd
3
4# setup
5wl0 = torch.tensor([700.0]) # nm
6k0 = 2 * torch.pi / wl0
7mat_particle = pmd.materials.MatDatabase("Si")
8n_env = 1.0
9
10# optimization
11r_init = 60.0 # initial guess
12r_opt = torch.tensor([r_init], requires_grad=True)
13optimizer = torch.optim.Adam(params=[r_opt], lr=0.5)
14
15# gradient loop
16for i in range(100):
17 optimizer.zero_grad()
18 particle = pmd.Particle(
19 mat_env=n_env,
20 r_layers=r_opt,
21 mat_layers=[mat_particle],
22 )
23 q_sca = particle.get_cross_sections(k0)["q_sca"]
24
25 loss = -q_sca # *maximize* scattering: minus sign
26 loss.backward() # calc. gradients via autodiff
27 optimizer.step() # apply gradients on params
28
29print("radius for strongest scattering:", r_opt.item())
Listing 3: Example how to use autodiff for a simple optimization.
Figure 5: “Mie-informed” tandem network. (a) Sketch of the training configuration. The design observable (e.g., scattering efficiency, or any other Mie-calculated value) is fed into the “inverse network” (here a simple MLP). The $N$ result is interpreted as core-shell geometry parameters and fed into PyMieDiff, which calculates the response of the particle analytically. The result of the Mie calculation is compared to the original input data using some loss function (here MSE). The inverse network is trained using backpropagation of the loss through the differentiable Mie solver and then the network itself. The inset shows the training loss for an inverse network trained on scattering efficiency spectra of dielectric particles. (b) Examples of inverse design responses (orange lines) vs. randomly sampled design targets (dashed black lines) from the training dataset. Spectral matching is done at $40$ points over the wavelength range between $400$ nm and $800$ nm.